OpenStudy (anonymous):

I tried solving this system of equations and got no solution. Could someone see if they get a solution? 8x-2y=5 -12x+3y=7

OpenStudy (anonymous):

(8x-2y=5)3 (-12x+3y=7)2 24x-6y=15 -24x+6y=14 15 = 14 There is no solution, the two lines are parallel.

OpenStudy (anonymous):

Ah yes! That's what I thought. I hate how math questions don't state "And if there's no solution, explain why." Instead they make you second guess your answer.

OpenStudy (anonymous):

If you go into advance math, most problems WON'T have a solution.

OpenStudy (anonymous):

Oh so than it becomes way more confusing. Yikes. Unfortunately, that's where I'm heading. Oh well.

OpenStudy (unklerhaukus):

If you rearrange both equations to the (slope-intercept) form y=mx+b you will find both have the same slope (m) but different intercepts (b) lines with the same slope are parallel , and hence can never intersect.

OpenStudy (unklerhaukus):

A solution to a system is/are the point(s) of intersection

OpenStudy (anonymous):

I see, I see. Thank you for explaining that in detail! :)

OpenStudy (unklerhaukus):

:•)